# IPPP/25/40

### v1

@@ Appendix A UCMH formation
In Fig. 3 we demonstrate how this affects our limits on $\mathcal{P}_{\mathcal{R}}(k)$, noting that CMB bounds in any case require collapse before recombination. The other constraints, from wide binary evaporation and lensing of the Icarus star, clearly show the generally expected behaviour outlined above: while the constraint on the power spectrum normalization tightens as a direct consequence of the decrease in $\delta_{c}$, the mass range where the constraints apply is reduced as the physical UCMH size increases. The actual reduction of the maximally constrained value of $k$ visible in the figure is further enhanced, compare to the naively expected $\propto(1+z_{c})$, by the fact that the Icarus bound itself is mass-dependent.
In reality, collapse would happen at all the redshifts indicated in the figure, so the overall constraint on $\mathcal{P}_{\mathcal{R}}(k)$ should in principle be taken as the envelope of the curves shown here. However, we re-iterate that these constraints become less and less robust as $z_{c}$ is decreased. For late-time collapse, a more realistic computation should thus take into account both survival probability and changes in the UCMH density profile due to merger events and other effects during non-linear structure formation. This must be based on full numerical simulations and is beyond the scope of this work.
@@
Figure 3: Dependence of the bounds on $\mathcal{P}_{\mathcal{R}}(k)$ for wide binary evaporation and Icarus constraints on the assumed redshift $z_{c}$ of collapse. CMB accretion bounds require formation before recombination.
Appendix B UCMH density profiles and growth by accretion
@@
As stressed in the main text, UCMHs forming very early and in isolation are expected to obtain very steep density profiles $\rho\propto r^{-\alpha}$ with $\alpha=9/4$. Once either of these assumptions is not met, however, somewhat less steep profiles with $\alpha\simeq 3/2$ are more realistic. Also independently of the UCMH context it is interesting to ask how our general EDO constraints are affected by the density profile for $r<r_{90}$. We show this in Fig. 4, demonstrating that the value of $\alpha$ only has a relatively mild impact on the limit. For both the wide binary evaporation constraint and the CMB accretion constraint, this remains true as long as the objects under consideration remain at least somewhat compact, implying that such limits are dominantly driven by the relation between $M$ and $R_{90}$ [35, 36]. The microlensing efficiency for extended objects displays some sensitivity to the density profile, as was discussed in Ref. [31], where the threshold impact parameter effectively encodes the efficiency (see in particular their Fig. 3). In contrast, the conservative assumption we adopt for applying the Icarus constraint assumes no such profile dependence.
@@
Figure 4: EDO bounds for different density profiles $\rho\propto r^{-\alpha}$ with fixed values of $R_{90}$ (blue) and the corresponding UCMH bounds (red). The solid lines are the same as in Fig. 1 in the main text.
Another aspect entering our limits is how we treat the UCMH mass. In the main text, we conservatively assume that $M_{\rm UCMH}$ is given by Eq. (1), with no subsequent growth due to accretion. While this behaviour corresponds to the finding in some simulations, other works have argued that $M_{\rm UCMH}$ should grow by a factor of $(1+z_{\rm eq})/(1+z_{c})$ during matter domination. We illustrate in Fig. 5 the impact of such a mass growth, concluding that the impact on the CMB accretion and Icarus limits on $\mathcal{P}_{\mathcal{R}}(k)$ is noticeable but still relatively small, while the impact on the wide binary evaporation limits is significant. In combination, this would imply that UCMH constraints outperform CMB spectral distortion constraints over the entire range of scales considered in our analysis. The reason why wide binary evaporation is particularly sensitive to the UCMH mass is simply that the constraint on $\mathcal{P}_{\mathcal{R}}(k)$ scales nonlinearly with the constraint on $f_{\rm DM}$, such that a fractional change in a less stringent constraint on $f_{\rm DM}$ affects $\mathcal{P}_{\mathcal{R}}(k)$ more strongly. We conclude our discussion of the potential impact of a UCMH mass growth by noting that a growth $\propto(1+z_{c})^{-1}$ can in any case not be the correct prescription for large values of $\mathcal{P}_{\mathcal{R}}(k)$: once there is a significant probability $\beta\gtrsim\mathcal{O}(10^{-3})$ for a region of a given size to collapse, cf. Eq. (3), there is no longer a sufficient amount of unbounded background matter in the surrounding regions to sustain such a mass growth (formally resulting in ‘overclosure constraints’ $\Omega_{\rm UCMH}>\Omega_{\rm DM}$ if applying the $\propto(1+z_{c})^{-1}$ scaling also in this regime).
@@
Appendix C Spectrum dependence of $\mu$-distortion and PTA constraints
@@
We stress that a fully scale-invariant power spectrum, $\mathcal{P}_{\mathcal{R}}(k)=const.$ for all $k$, is largely an academic exercise that we only include here for illustration: for normalizations larger than $\sim 10^{-9}$, both CMB spectral distortion and PTA constraints become irrelevant as such a choice is completely excluded by large-scale CMB temperature anisotropy observations. For direct comparison with the literature [28], we also include a (blue) line corresponding to the range of scales $0.15\lesssim k/k_{R}\lesssim 12.35$ that contribute $99$% to the mass variance $\sigma_{\chi,H}^{2}$ given in Eq. (4).
@@
In Fig. 2 of the main text, we have presented results for a spectrum that is locally scale-invariant, implicitly assuming that $\mathcal{P}_{\mathcal{R}}(k)$ is non-vanishing only for a relatively small window in $k$ because the UCMH constraints are anyway only sensitive to scales $k\sim k_{R}$. Fig. 6 demonstrates that the UCMH constraints will clearly dominate over both CMB spectral distortion and PTA constraints even for much ‘broader’ power spectra, with $\mathcal{P}_{\mathcal{R}}(k)\simeq const.$ over a comparatively large range of scales.
Appendix D Constraints from Highly Magnified Stars
Observations of individual background stars near cluster caustics have opened a novel avenue for constraining compact dark matter. The smooth mass distribution of a galaxy cluster by itself produces a critical curve in the image plane and a corresponding fold caustic in the source plane whenever the projected surface density approaches the critical value. A secondary, smaller-scale lens, such as an intracluster (ICL) star or compact object?perturbs this macro caustic, creating a web of microcaustics around it. In favourable configurations, where a background star lies close to this combined macro?micro caustic, the magnification can reach extreme values ($\mu\sim 10^{3}$ or more), but the detailed light curve depends sensitively on whether the caustic remains smooth or is fragmented by compact lenses.
@@ Appendix D Constraints from Highly Magnified Stars
Figure 6: Comparison of the bounds from CMB $\mu$ distortions and PTAs for different integration ranges over a scale-invariant spectrum, as indicated. The blue line corresponds to a range $0.15\lesssim k/k_{R}\lesssim 12.35$, following Ref. [28].
@@
As described in the main text, the first such constraint was derived from “Icarus” (MACS J1149 LS1 at $z=1.49$), whose magnification reached $\mu\sim 2000$  [71, 53, 54]. For compact lenses, three mass regimes are imporant: (i) for small lenses, the effective Einstein radius ($\bar{R}_{E}\propto\sqrt{M_{\rm EDO}}$) is smaller than the stellar radius, so the star smooths over microcaustics and no constraint arises; (ii) for intermediate masses, the magnification is reduced in a way that depends only on the total surface mass density in compact lenses, yielding a constraint on $f_{\rm DM}$ independent of $M$; (iii) for large $M$, the EDO itself acts as the main lens, and requiring a detectable caustic crossing implies $f_{\rm DM}\propto M^{1/3}$, with no constraint for $M\gtrsim 10^{4}M_{\odot}$.
@@
$\mu_{t,\rm max}=\frac{M}{\pi\Sigma R_{E}^{2}}\approx 1.2f_{\rm DM}^{-1}\,,$ (6)
@@
Crucially, the derivation of the Icarus constraint in [53] on point-like lenses assumed point sources, an assumption which breaks down when $\bar{R}_{E}\lesssim R_{\star}$, as microcaustics are then averaged over by the source’s finite size. A more realistic treatment requires ray-tracing simulations with extended sources and finite-size lenses (e.g. [54]). In part of the regime, complementary constraints were derived based in the expected event rate [74].
Additional events promise to refine or strengthen these constraints. The “Godzilla” event in the Sunburst Arc ($z=2.37$) reached $\mu\sim 600$ with microcaustic peaks up to a few $10^{3}$ [75]; “Spock” transients in MACS J0416 showed $\mu\sim 400$ [76]; “Mothra” may exceed $\mu\sim 6000$ if its compact perturber is confirmed [77]. “Earendel” (at $z=6.2$) which has been found to have $\mu>4000$ from JWST imaging [78]. [^2] However, lens models in Ref. [79] suggest a lower value of $\mu\approx 43$-$67$. This discrepancy highlights the sensitivity of such constraints to lens modeling systematics and underscores the need for continued development of cluster mass reconstructions. These increasingly numerous caustic-crossing stars may offer a pathway to probe smaller compact objects than previously accessible. Since $\bar{R}_{E}\propto\sqrt{\mu_{t}}$, higher-magnification systems enhance sensitivity to lower-mass, as well as more extended, lenses. Future constraints will likely benefit from improved ray-tracing and lens modeling, as well as a statistical sample of such events across multiple clusters.

### v2

@@ Appendix A Appendix A: UCMH formation
In Fig. 3 we demonstrate how our limits on $\mathcal{P}_{\mathcal{R}}(k)$ depend on the assumed value of $z_{c}$, noting that CMB bounds in any case require collapse before recombination. The other constraints, from wide binary evaporation and lensing of the Icarus star, clearly show the generally expected behaviour outlined above: while the constraint on the power spectrum normalization tightens as a direct consequence of the decrease in $\delta_{c}$, the mass range where the constraints apply is reduced as the physical UCMH size increases. The actual reduction of the maximally constrained value of $k$ visible in the figure is stronger than the naively expected $\propto(1+z_{c})$, because the Icarus bound itself is mass-dependent.
@@
Figure 3: Wide binary evaporation and Icarus constraints on $\mathcal{P}_{\mathcal{R}}(k)$, for various assumed redshifts $z_{c}$ of collapse. CMB accretion bounds require formation before recombination.
In reality, collapse would happen at all the redshifts indicated in the figure, so the overall constraint on $\mathcal{P}_{\mathcal{R}}(k)$ should in principle be taken as the envelope of the curves shown here. However, we re-iterate that these constraints become less and less robust as $z_{c}$ is decreased. For late-time collapse, a more realistic computation should thus take into account both survival probability and changes in the UCMH density profile due to merger events and other effects during non-linear structure formation. This must be based on full numerical simulations and is beyond the scope of this work. Taking into account earlier collapse times, as also shown in Fig. 3, is particularly relevant for the Icarus constraint. This is because the microlenses must be smaller than their Einstein radius, $1>R^{(90)}_{\rm UCMH}/R_{E}\propto k^{\frac{1}{2}}(1+z_{c})$, such that a larger $z_{c}$ allows to probe larger $k$. At the same time, $R_{E}\propto\sqrt{M_{\rm UCMH}}$ must remain larger than the stellar radius (see also the discussion further down), corresponding to the $z_{c}$-independent bound $k\lesssim 2\cdot 10^{5}\,{\rm Mpc}^{-1}$.
Appendix B Appendix B: UCMH density profiles and growth by accretion
@@
As stressed in the main text, UCMHs forming very early and in isolation are expected to obtain very steep density profiles $\rho\propto r^{-\alpha}$ with $\alpha=9/4$. Once either of these assumptions is not met, however, somewhat less steep profiles with $\alpha\simeq 3/2$ are more realistic. Also independently of the UCMH context it is interesting to ask how our general EDO constraints are affected by the density profile for $r<r_{90}$. We show this in Fig. 4, demonstrating that the value of $\alpha$ only has a relatively mild impact on the limits. For both the wide binary evaporation constraint and the CMB accretion constraint, this remains true as long as the objects under consideration remain at least somewhat compact, implying that such limits are dominantly driven by the relation between $M$ and $R_{90}$ [39, 40]. The microlensing efficiency for extended objects displays some sensitivity to the density profile, as was discussed in Ref. [35], where the threshold impact parameter effectively encodes the efficiency (see in particular their Fig. 3). In contrast, the conservative assumption we adopt for applying the Icarus constraint assumes no such profile dependence.
@@
Figure 4: EDO bounds for different density profiles $\rho\propto r^{-\alpha}$, for fixed $R_{90}$ (blue) and for UCMHs (red), respectively. The solid lines are the same as in Fig. 1 in the main text.
Another aspect entering our limits is how we treat the UCMH mass. In the main text, we conservatively assume that $M_{\rm UCMH}$ is given by Eq. (1), with no subsequent growth due to accretion. While this behaviour corresponds to the finding in some simulations, other works have argued that $M_{\rm UCMH}$ should grow by a factor of $(1+z_{\rm eq})/(1+z_{c})$ during matter domination. We illustrate in Fig. 5 the impact of such a mass growth, concluding that the impact on the CMB accretion and Icarus limits on $\mathcal{P}_{\mathcal{R}}(k)$ is noticeable but still relatively small, while the impact on the wide binary evaporation limits is significant. In combination, this would imply that UCMH constraints outperform CMB spectral distortion constraints over the entire range of scales considered in our analysis. The reason why wide binary evaporation is particularly sensitive to the UCMH mass is simply that the constraint on $\mathcal{P}_{\mathcal{R}}(k)$ scales nonlinearly with the constraint on $f_{\rm DM}$, such that a fractional change in a less stringent constraint on $f_{\rm DM}$ affects $\mathcal{P}_{\mathcal{R}}(k)$ more strongly. We conclude our discussion of the potential impact of a UCMH mass growth by noting that a growth $\propto(1+z_{c})^{-1}$ can in any case not be the correct prescription for large values of $\mathcal{P}_{\mathcal{R}}(k)$: once there is a significant probability $\beta\gtrsim\mathcal{O}(10^{-3})$ for a region of a given size to collapse, cf. Eq. (3), there is no longer a sufficient amount of unbounded background material in the surrounding regions to sustain such a mass growth (formally resulting in ‘overclosure constraints’ $\Omega_{\rm UCMH}>\Omega_{\rm DM}$ if applying the $\propto(1+z_{c})^{-1}$ scaling also in this regime).
@@
Appendix C Appendix C: $\mu$-distortion and PTA constraints
@@
We stress that a fully scale-invariant power spectrum, $\mathcal{P}_{\mathcal{R}}(k)=const.$ for all $k$, is largely an academic exercise that we only include here for illustration: for normalizations larger than $\sim 10^{-9}$, both CMB spectral distortion and PTA constraints become irrelevant as such a choice is completely excluded by large-scale CMB temperature anisotropy observations. For direct comparison with the literature [32], we also include a (blue) line corresponding to the range of scales $0.15\lesssim k/k_{R}\lesssim 12.35$ that contribute 9999% to the mass variance $\sigma_{\chi,H}^{2}$ given in Eq. (4).
@@
In Fig. 2 of the main text, we have presented results for a spectrum that is locally scale-invariant, implicitly assuming that $\mathcal{P}_{\mathcal{R}}(k)$ is non-vanishing only for a relatively small window in $k$ because the UCMH constraints are anyway only sensitive to scales $k\sim k_{R}$. Fig. 6 demonstrates that the UCMH constraints (shown in Fig. 2 in the main text) will clearly dominate over both CMB spectral distortion and PTA constraints even for much ‘broader’ power spectra, with $\mathcal{P}_{\mathcal{R}}(k)\simeq const.$ over a comparatively large range of scales.
@@
Figure 6: Comparison of the bounds from CMB $\mu$ distortions and PTAs for different integration ranges over a scale-invariant spectrum, as indicated. The purple line corresponds to what we adopted in the main text, while the blue line corresponds to a range $0.15\lesssim k/k_{R}\lesssim 12.35$, following Ref. [32].
Appendix D Appendix D: Constraints from Highly Magnified Stars
Observations of individual background stars near cluster caustics have opened a novel avenue for constraining compact DM. The smooth mass distribution of a galaxy cluster by itself produces a critical curve in the lensed image plane and a corresponding fold caustic in the source plane whenever the projected surface density approaches the critical value. A secondary, smaller-scale lens, such as an intracluster (ICL) star or compact object perturbs this macro caustic, creating a web of microcaustics around it. In favourable configurations, where a background star lies close to this combined macro-micro caustic, the magnification can reach extreme values, $\mu\sim 10^{3}$ or more, but the light curve depends sensitively on whether the caustic remains smooth or is fragmented by compact lenses.
@@ Appendix D Appendix D: Constraints from Highly Magnified Stars
The first such constraint was derived from “Icarus” (MACS J1149 LS1 at $z=1.49$), whose magnification reached $\mu\sim 2000$ [80, 62, 63] (with macro-lens magnification $\mu_{t}\sim 100$). For compact microlenses, $\bar{R}_{E}<R_{90}$, three mass regimes are important: (i) for small lenses, the effective Einstein radius ($\bar{R}_{E}\propto\sqrt{M_{\rm EDO}}$) is smaller than the stellar radius, so the star smooths over microcaustics and no constraint arises; (ii) for intermediate masses, the magnification is reduced in a way that depends only on the total surface mass density in (all) compact lenses, yielding a constraint on $f_{\rm DM}$ independent of $M$; (iii) for large $M$, the EDO itself acts as the main lens, and requiring a detectable caustic crossing implies $f_{\rm DM}\propto M^{1/3}$, with no constraint for $M\gtrsim 10^{4}M_{\odot}$.
@@
$\mu_{t,\rm max}=\frac{M}{\pi\Sigma R_{E}^{2}}\approx 1.2f_{\rm DM}^{-1}\,$,\\ (6)
@@
Crucially, the derivation of the Icarus constraint in Ref. [62] on point-like lenses assumed point sources, an assumption which breaks down when $\bar{R}_{E}\lesssim R_{\star}$, as microcaustics are then averaged over by the source’s finite size. A more realistic treatment requires ray-tracing simulations with extended sources and finite-size lenses (e.g. Ref. [63]). In part of the regime, complementary constraints were derived based on the expected event rate [83].
Additional events promise to refine or strengthen these constraints. The “Godzilla” event in the Sunburst Arc ($z=2.37$) reached $\mu\sim 600$ with microcaustic peaks up to a few $10^{3}$ [84]; “Spock” transients in MACS J0416 showed $\mu\sim 400$ [85]; “Mothra” may exceed $\mu\sim 6000$ if its compact perturber is confirmed [86]. “Earendel” (at $z=6.2$) has been found to have $\mu>4000$ from JWST imaging [87]. [^2] However, lens models in Ref. [88] suggest a lower value of $\mu\approx 43$-6767. This discrepancy highlights the sensitivity of such constraints to lens modeling systematics and underscores the need for continued development of cluster mass reconstructions. These increasingly numerous caustic-crossing stars may offer a pathway to probe smaller compact objects than previously accessible. Since $\bar{R}_{E}\propto\sqrt{\mu_{t}}$, higher-magnification systems enhance sensitivity to lower-mass, as well as more extended, lenses. Future constraints will likely benefit from improved ray-tracing and lens modeling, as well as a statistical sample of such events across multiple clusters.